OpenStudy (anonymous):

Factor the trinomial y^3-15y^2+54y

OpenStudy (vishweshshrimali5):

Is it 15^2 ????

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sorry 15y^2

OpenStudy (vishweshshrimali5):

take y as common

OpenStudy (anonymous):

y^3-15y^2+54y =y(y^2-15y+54) =y(y-6)(y-9)

OpenStudy (vishweshshrimali5):

now factorise it

OpenStudy (vishweshshrimali5):

\[\large y(y(y-6)-9(y-6))\] \[\large y(y-9)(y-6)\]

OpenStudy (vishweshshrimali5):

got that @jlmack07

OpenStudy (anonymous):

I do and thank you very much for all your help

OpenStudy (vishweshshrimali5):

Never mind.